OpenStudy (destinykiara99):

Find the inverse of the function. f(x) = 3x - 2

OpenStudy (destinykiara99):

Answer Choices: f-1(x) = Quantity x plus two divided by three. f-1(x) = x divided by three. + 2 f-1(x) = Quantity x minus two divided by three. Not a one-to-one function

OpenStudy (rippedblue42):

To find the inverse, make f(x) y and switch y and x then make y f(x) again

OpenStudy (rippedblue42):

If you need me to check you answer I will :)

OpenStudy (destinykiara99):

I got A, is that correct? :)

OpenStudy (rippedblue42):

Do you need more help?

OpenStudy (rippedblue42):

Yes it is;)

OpenStudy (destinykiara99):

Yeah, I have another question, too. If you don't mind helping!

OpenStudy (rippedblue42):

No problem:)

OpenStudy (destinykiara99):

Describe how to transform the graph of f into the graph of g. f(x) = x4 and g(x) = -x4

OpenStudy (rippedblue42):

Ask/ metal/ fan awayXD

OpenStudy (rippedblue42):

What happens if a graph is negative?

OpenStudy (destinykiara99):

Answer Choices: Reflect the graph of f across the x-axis. Shift the graph of f down 1 unit. Reflect the graph of f across the x-axis and then reflect across the y-axis. Reflect the graph of f across the y-axis.

OpenStudy (rippedblue42):

Well actually first think about a regular positive graph. It goes up right? So the opposite of positive and negative and the opposite of up is down to negative graphs go down and practice up so one graph goes up and one down

OpenStudy (destinykiara99):

I'm thinking that it would be C out of my answer choices? @rippedblue42

OpenStudy (rippedblue42):

C is correct!

OpenStudy (destinykiara99):

Yay!

OpenStudy (rippedblue42):

Fan and metal me if I was helpful:)

OpenStudy (destinykiara99):

Done! Thank you!